### v1

# Ab Initio Framework of Electron-Phonon Coupling for Chiral Phonons with Giant Phonon Magnetic Moments in Magnetic Materials

Abstract
Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, this field remains severely limited due to the inability to accurately account for electron-phonon-coupling (EPC) effects for chiral phonons. Here, we resolve this challenge through developing an ab initio framework for computing EPC-driven phonon magnetic properties, applicable to all insulating and metallic materials. Our theory predicts a dramatic enhancement of phonon magnetic moments exceeding classical predictions by multiple orders, accompanied by distinct Zeeman splitting. Interestingly, these splittings induce observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states. Through first-principles calculations, we identify five topological phonon magnets with intrinsic phonon Chern states and also propose to detect low-energy neutral particles such as dark matters. This research not only establishes a computational framework for investigating EPC-driven phonon magnetic properties, but also provides a material design strategy for realizing topological phonon states in quantum magnetic materials.
Introduction. Chiral phonons, defined as collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and his cooperators [1, 2] and subsequently verified experimentally in monolayer WSe2 [3]. These chiral vibrations not only reveal fundamental properties of phonons previously unrecognized, but also establish a novel degree of freedom for manipulating phonon-mediated phenomena [1, 4, 5, 6, 7]. Recently, there has been a surge of research interest in chiral phonons, spanning theoretical investigations [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] to experimental explorations [19, 20, 21, 22, 23, 24, 25, 26, 27]. Of particular importance are phonon magnetic moments (PMMs) and their associated phonon Zeeman splittings, which constitute key mechanisms in chiral-phonon-involved effects [6, 7]. However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below practical application requirements [13], posing significant limitations for the advancing development and application of chiral phonons. Although electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [11, 13, 16], a unified ab initio theoretical framework of EPC-driven PMMs for realistic materials is lacking.
Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-driven PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions, which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons. The classical Zeeman splitting of chiral phonons is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger. This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
Besides magnetic moments, phonons can also demonstrate the Hall effect [28, 29, 30, 31] and manifest topological characteristics, such as Berry curvature and Chern numbers [32, 33, 34]. When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, essentially realizing phonon Chern insulators. These topological phonon states featuring topologically protected edge currents could significantly enhance both detection sensitivity and practical exploitation of the phonon Hall effect. Nevertheless, conventional TRS-breaking approaches for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations or induce parasitic effects [35, 36, 37]. Consequently, phonon Chern states driven by intrinsic magnetic moments without external fields are highly desirable. Crucially, predicting such intrinsic topological phonon states in magnets demands precise calculations of TRS-broken phonon spectra, which remains challenging due to the lack of ab initio framework for magnetic phonon systems [38, 39, 40, 41, 42, 43].
To resolve these challenges, we establish a unified ab initio framework for determining EPC-driven phonon magnetic properties, which enables first-principles calculations of the previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals. Our theory demonstrates EPC-driven enhancements of PMMs by four orders of magnitude, inducing giant chiral phonon Zeeman splittings. Building on these giant splittings, we propose the inertially decoupled model to realize intrinsic phonon Chern states in magnets. Furthermore, through first-principles calculations, we identify five promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$. Our work not only paves a material-design pathway for topological phonon magnets, but also establishes the framework for chiral phonon physics in magnetic phenomena.
Ab initio framework for phonon Zeeman energy. Previous theories regarding the magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals and are only effective for specific high-symmetry phonon wave vectors [11, 13, 14, 16]. However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop a unified theory that can comprehensively describe phonon behaviors in insulators, semimetals, and metals.
We begin with a TRS-breaking perturbation phonon Hamiltonian, $\delta H=-\sum_{ab}\eta_{ab}p_{a}u_{b}$, known as the Raman spin-phonon interaction [44], where $\eta_{ab}$ is a real antisymmetric matrix that satisfies $\eta_{ab}=-\eta_{ba}$. Here, $u_{a}=\sqrt{m}x_{a}$ represents the renormalized atomic displacement coordinate and $p_{a}$ represents the canonical momentum related to $u_{a}$, where $m$ denotes the atomic mass. By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$. Here, we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$. Therefore, the Zeeman energy of the chiral phonon can be regarded as an effective linear response to its angular momentum.
We notice that $\bm{\eta}$ can be deduced from the TRS-breaking self-energy correction of chiral phonons driven by EPC. Remarkably, within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig. 1(b)(see the supplement material (SM)) [45],
$\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (1)
Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively. $N$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$. $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [46, 47]. Notably, all parameters in Eq. (1) can be directly derived from first-principles calculations, obviating the need for empirical inputs or additional assumptions. As a result, the previously inaccessible TRS-breaking phonon spectra driven by EPC in both insulating and metallic systems is now accessible across the whole Brillouin zone.
Phonon Zeeman splitting and enhanced phonon $g$-factor. We propose the essential conditions for realizing significant phonon Zeeman splittings by EPC. Since the primary determinant of the phonon Zeeman splitting is the susceptibility $\bm{\chi}^{-}$, there are two key points according to Eq. (1): (i) The contribution of EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is substantial, and materials composed of light atoms tend to possess strong $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$. (ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, the parametric sensitivity in Eq. (1) leads to an enhancement of $\bm{\chi}^{-}$ [45]. Consequently, we can conclude that EPC-driven large phonon Zeeman splittings are most likely to occur in narrow-gap semiconductors, semimetals and metals comprising light elements.
Large phonon Zeeman splittings suggest that phonons exhibit significant phonon magnetic moments. To derive the PMM, we taking the derivative $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}$ and the phonon Zeeman energy is further rewritten as $E_{\text{Z}}=-\bm{B}\cdot\bm{\mu}^{\text{ph}}$, where $\bm{B}$ represents the effective magnetic field, and $\bm{\mu}^{\text{ph}}$ denotes the PMM (see SM  [45]),
$\mu^{\text{ph}}_{c}=-g^{\text{ph}}_{c}\frac{\mu_{B}}{\hbar}s^{\text{ph}}_{c}.$ (2)
Here, $\mu_{B}$ denotes the Bohr magneton, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon, where $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$, $g^{e}_{c}$ is the $g$-factor of electrons along $c$ direction involved in EPC, and the dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor. This amplification factor is primarily influenced by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electronic states, and it can significantly exceed unity. This enhancement allows the $g$-factor of the chiral phonon to the order of $g^{e}_{c}$.
This EPC-enhanced PMM can be intuitively understood through the following mechanism. As chiral phonons drive atomic cyclotron motion, a portion of its angular momentum is transferred to the electrons via EPC. Owing to the considerably small mass of electrons, the angular momentum transfer gives rise to strong loop electron currents, thereby resulting in a significant PMM. In contrast, within the classical framework, the phonon’s $g$-factor is constrained to the order of $10^{-5}\sim 10^{-4}g^{e}_{c}$ due to heavy ionic masses compared to the electron mass [13, 48].
Figure 2: Inertially decoupled model (IDM) and intrinsic phonon Chern states. (a) Schematic of optical chiral phonon modes: When $m\sim M$, both atoms contribute to the high-frequency model (up); in the IDM condition ($m\ll M$), the optical model is primarily contributed by light atoms (down). (b) Phonon dispersions of nonmagnetic honeycomb lattice and magnetic triangle lattice. (c-e) TRS-breaking phonon dispersions of 2D/3D IDM structures: the 2D IDM by combining the honeycomb and triangular lattices (c,d), and the 3D IDM by stacking 2D IDM along the $z$-direction (e). The color gradient from violet and yellow represents the contributions from the honeycomb and triangular lattice, and the colored gradient from blue and red represents the phonon angular momentum magnitude. (f) Edge states (cyan/magenta) of IDM under open boundaries. Parameters are detailed in the SM [45].
Intrinsic phonon Chern states and the inertially decoupled model. The large phonon Zeeman splitting can pave the way toward a phonon Chern state through the phonon band inversion [45]. Following the analogy to the Chern state in the Haldane model [49], we adopt the two-dimensional (2D) honeycomb lattice [44, 50, 51] as an illustrative example. When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $m_{I}$ and $m_{T}$, respectively. It has been shown that when $|m_{T}|>|m_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [44].
However, realizing such phonon Chern states in realistic materials hosting honeycomb lattices remains challenging. Although Coriolis forces and strong magnetic fields can produce a nonzero $m_{T}$ [44], these approaches either require large external fields or complex experimental measurements. In contrast, the EPC-enhanced phonon Zeeman splitting could offer a sizable $m_{T}$ in magnets without the need of external fields. To further satisfy the condition $|m_{T}|>|m_{I}|$, it is necessary to minimize $|m_{I}|$. Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|m_{I}|=0$ would be the ideal candidate. Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered. Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $m_{I}$, which will make the condition $|m_{T}|>|m_{I}|$ difficult to satisfy.
To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference. A prototypical example of this model is constructed by combining a honeycomb lattice composed of nonmagnetic light atoms of mass $m$ with a triangular lattice of heavy magnetic atoms of mass $M$ ($m\ll M$), as shown in Figs. 2(b,c). Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions mainly originate from the light and heavy atoms, respectively, as shown in Figs. 2(a,c), which is termed “inertially decoupled” (see SM [45] for details). Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $m_{I}$ mass at both the $K$ and $K^{\prime}$ valleys. Meanwhile, the magnetic $M$ atoms breaks TRS and provide effective magnetic fields to generate large Zeeman splittings of chiral phonons via EPC, producing the desired large $m_{T}$ mass. The large $m_{T}$ opens a topologically nontrivial phonon gap (seen in Fig. 2(c,d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by the topological edge states crossing the phonon gap, as shown in Fig. 2(f).
The inertially decoupled model is also applicable to three-dimensional (3D) magnets. Interestingly, when the nonmagnetic honeycomb lattice (the $m$ sublayer) and the magnetic triangular lattice (the $M$ sublayer) are stacked alternately along the $z$-direction, each $M$ sublayer effectively acts as a buffer layer that nearly decouples the phonon dispersions of adjacent $m$ sublayers in the high-frequency regime. A representative phonon dispersion of such a 3D bulk model is shown in Fig. 2(e). Notably, since each $m$ sublayer realizes a Chern state with $|C|=1$, the system exhibits an intrinsic 3D phonon Chern state. This configuration hosts multiple chiral phonon edge states resulting from a total Chern number $|C|=N_{z}$, where $N_{z}$ is the number of $m$ sublayers.
Figure 3: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$. The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$. (c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the TRS breaking, respectively. (e) The phonon dispersion of $\mathrm{EuSi_{2}}$ under open-boundary conditions. (f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the TRS breaking, respectively. (h) The phonon dispersion of the sandwiched slab under open-boundary conditions. The color gradient from red to blue represents the magnitude of the phonon angular momentum. Cyan and magenta label the surface or edge modes originating from opposite boundaries, respectively.
Candidate materials. Building upon the above theoretical framework, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [46, 47] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from dynamical matrix through near-degenerate perturbation theory (see SM [45]), yielding TRS-broken phonon dispersions.
Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal $P6/mmm$ magnets, specifically $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [58], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [59], and $\mathrm{CrB_{2}}$ [60, 61], exhibiting intrinsic phonon Chern states via EPC-enhanced Zeeman splittings ($\Delta_{\text{z}}=2|E_{\text{Z}}|$). As summarized in Table I, they all have large observable phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\gamma$ is phonon linewidth resulting from the EPC [62, 63]. Note that the Zeeman splitting should be larger than the phonon linewidth to produce an observable phonon gap.
Here, we take $\mathrm{EuSi_{2}}$ as a representative example (see SM [45] for details on all candidate materials). The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs. 3(a, e). Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a magnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction. The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately $28\,u$ and $152\,u$ ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$. As a result, $\mathrm{EuSi_{2}}$ can be aptly described by our proposed inertially decoupled model.
Our first-principles calculations reveal that the EPC induces a pronounced Zeeman splitting of chiral phonons in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs. 3(c, d)], opening a sizable full gap ($\approx 0.35$ THz) throughout the Brillouin zone. Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state. This is further supported by the presence of boundary modes in the phonon dispersions of $\mathrm{EuSi_{2}}$ under open-boundary conditions, as illustrated in Fig. 3(e). To achieve a 2D Chern state, we construct a sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig. 3(b). By comparing Figs. 2(f) and 2(g), we can see that the Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ in the phonon dispersion of this sandwiched slab, and the accompanying single pair of edge modes traversing the phonon gap, as shown in Fig. 3(h), provide definitive evidence for intrinsic phonon Chern states.
Table 1: The direct phonon gaps at K points for the five candidate materials, where $\gamma$ and $\Delta_{\text{z}}$ represent the phonon linewidth and the phonon Zeeman splitting, respectively. The observable phonon gap $\Delta_{\text{ph}}$ is defined as $\Delta_{\text{z}}-\gamma$.
Mat. $\Delta_{\text{z}}$ (THz) $\gamma$ (THz) $\Delta_{\text{ph}}$ (THz)
$\mathrm{EuSi_{2}}$ 0.426 0.049 0.377
$\mathrm{GdSi_{2}}$ 0.406 0.061 0.345
$\mathrm{EuGa_{2}}$ 0.134 0.028 0.106
$\mathrm{GdGa_{2}}$ 0.145 0.015 0.130
$\mathrm{CrB_{2}}$ 0.270 0.033 0.237
Discussion and conclusion. The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring extreme external strong magnetic fields, Coriolis forces, or stress gradients. One potential advantage may be its ability to detect neutral particles by converting them into phonon edge currents. Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging. However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within that of the first “conduction” band above the topological gap in the phonon dispersion, this process may potentially excite the corresponding phonon modes [64, 65], driving the system out of equilibrium. As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via edge states, producing a edge phonon current immune to scattering[Fig. 4(a)]. This robust edge current is significantly more measurable than conventional phonon excitations dispersed throughout the material. However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents. Thus, the gap determines the maximum operating temperature for such a detector. Fig. 4(b) lists the transfered energy ranges and gaps for all five candidate materials (see SM [45] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$. This energy range corresponds to light-mass dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates.
Figure 4: Detectors of neutral particles based on the intrinsic phonon Chern state. (a) Schematic illustrations of the detection process. Neutral particles excite phonon modes in high-frequency “conduction” bands. These modes then gradually relax to lower-frequency regimes via chiral edge phonon channels. (b) Transfered energy ranges and phonon Chern gaps for all the layered candidate materials.
In summary, we have established a unified ab initio framework enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems. Under this framework, we propose an inertially decoupled model to design intrinsic phonon Chern states, and identify five candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz. Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through phonon edge current transduction. These advances simultaneously open broad pathways for revealing new phonon magnetic effects in quantum magnetic materials.
Acknowledgements.
The authors thank Lei Zhang and Zuowei Liu for valuable discussions. This work is partly supported by National Key Projects for Research and Development of China (Grants No. 2024YFA1409100, No. 2021YFA1400400 and No. 2022YFA1403602), the Natural Science Foundation of Jiangsu Province (No. BK20220032, BK20233001, BK20243011), the Natural Science Foundation of China (No. 12074181, No. 12104217, No. 12174182, and No. 92365203), and the e-Science Center of Collaborative Innovation Center of Advanced Microstructures.

### v2

@@
-Ab Initio Framework of Electron-Phonon Coupling for Chiral Phonons with Giant Phonon Magnetic Moments in Magnetic Materials
+Ab Initio theory of Electron-phonon-coupling-induced Giant Magnetic Moments of Chiral Phonons in Magnetic Materials
 Abstract
 Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention.
-However, this field remains severely limited due to the inability to accurately account for electron-phonon-coupling (EPC) effects for chiral phonons.
-Here, we resolve this challenge through developing an ab initio framework for computing EPC-driven phonon magnetic properties, applicable to all insulating and metallic materials.
-Our theory predicts a dramatic enhancement of phonon magnetic moments exceeding classical predictions by multiple orders, accompanied by distinct Zeeman splitting.
-Interestingly, these splittings induce observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states.
-Through first-principles calculations, we identify five topological phonon magnets with intrinsic phonon Chern states and also propose to detect low-energy neutral particles such as dark matters.
-This research not only establishes a computational framework for investigating EPC-driven phonon magnetic properties, but also provides a material design strategy for realizing topological phonon states in quantum magnetic materials.
+However, a long-standing critical issue in this field is the lack of ab initio method to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC).
+Here, we resolve this challenge by developing an ab initio theory for calculating EPC-induced phonon magnetic properties, applicable to both insulating and metallic materials.
+Based on this theory, we demonstrate EPC-induced giant phonon magnetic moments and resulting phonon Zeeman splittings in magnetic metals, which are orders of magnitude larger than classical predictions from point-charge models.
+Interestingly, these splittings could open observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states.
+Through first-principles calculations, we first propose candidate materials with such intrinsic phonon Chern states hosting robust edge phonon currents which may be applied to detecting neutral particles such as dark matter particles.
+Our work not only establishes a theoretical foundation for EPC-induced phonon magnetic properties, but also enables the ab initio calculation of long-sought TRS-breaking phonon spectra throughout the Brillouin zone in realistic materials.
 Introduction.
-Chiral phonons, defined as collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and his cooperators [1, 2] and subsequently verified experimentally in monolayer WSe2 [3].
-These chiral vibrations not only reveal fundamental properties of phonons previously unrecognized, but also establish a novel degree of freedom for manipulating phonon-mediated phenomena [1, 4, 5, 6, 7].
-Recently, there has been a surge of research interest in chiral phonons, spanning theoretical investigations [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] to experimental explorations [19, 20, 21, 22, 23, 24, 25, 26, 27].
-Of particular importance are phonon magnetic moments (PMMs) and their associated phonon Zeeman splittings, which constitute key mechanisms in chiral-phonon-involved effects [6, 7].
-However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below practical application requirements [13], posing significant limitations for the advancing development and application of chiral phonons.
-Although electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [11, 13, 16], a unified ab initio theoretical framework of EPC-driven PMMs for realistic materials is lacking.
-Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-driven PMM.
+Chiral phonons, collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and Niu [1, 2] and subsequently verified experimentally in monolayer WSe2 [3].
+Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and experiments [15, 16, 17, 18, 19, 20, 21, 22, 23].
+Of particular importance are phonon magnetic moments (PMMs) of chiral phonons and their associated phonon Zeeman splittings, playing a crucial role in chiral-phonon-involved effects [1, 24, 25, 26, 27].
+However, classical models based on point-charge approximations yield extremely small PMM values due to the significant large mass of nucleus  [9].
+While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [7, 9, 12], an ab initio method capable of accurately calculating EPC-induced PMMs in general materials is still lacking.
+Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-induced PMM.
 In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions, which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons.
-The classical Zeeman splitting of chiral phonons is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger.
+The classical Zeeman splitting of chiral phonons deduced from the point-charge model is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger.
 This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
-Besides magnetic moments, phonons can also demonstrate the Hall effect [28, 29, 30, 31] and manifest topological characteristics, such as Berry curvature and Chern numbers [32, 33, 34].
-When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, essentially realizing phonon Chern insulators.
-These topological phonon states featuring topologically protected edge currents could significantly enhance both detection sensitivity and practical exploitation of the phonon Hall effect.
-Nevertheless, conventional TRS-breaking approaches for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations or induce parasitic effects [35, 36, 37].
-Consequently, phonon Chern states driven by intrinsic magnetic moments without external fields are highly desirable.
-Crucially, predicting such intrinsic topological phonon states in magnets demands precise calculations of TRS-broken phonon spectra, which remains challenging due to the lack of ab initio framework for magnetic phonon systems [38, 39, 40, 41, 42, 43].
-To resolve these challenges, we establish a unified ab initio framework for determining EPC-driven phonon magnetic properties, which enables first-principles calculations of the previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals.
-Our theory demonstrates EPC-driven enhancements of PMMs by four orders of magnitude, inducing giant chiral phonon Zeeman splittings.
-Building on these giant splittings, we propose the inertially decoupled model to realize intrinsic phonon Chern states in magnets.
-Furthermore, through first-principles calculations, we identify five promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$.
-Our work not only paves a material-design pathway for topological phonon magnets, but also establishes the framework for chiral phonon physics in magnetic phenomena.
-Ab initio framework for phonon Zeeman energy.
+Besides magnetic moments, phonons can also exhibit Hall effects [28, 29, 30, 31] and topological characteristics, such as Berry curvature and Chern numbers [32, 33, 34].
+When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents with significantly enhanced detectability compared to conventional phonon Hall effects.
+Nevertheless, TRS-breaking external fields for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations [35, 36].
+Consequently, phonon Chern states driven by the phonon Zeeman splitting in magnets without external fields are highly desirable.
+However, predicting such intrinsic topological phonon states in magnets remains challenging due to the inability to perform ab initio calcualtions of TRS-broken phonon spectra [37, 38, 39, 40, 41, 42].
+To resolve these challenges, we establish an ab initio theory for determining EPC-induced phonon magnetic properties, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals.
+Our theory relies on treating the angular momentum of chiral phonon as an effective perturbative field to the electron-phonon coupled system, with the phonon Zeeman energy interpreted as the response to this perturbation.
+Based on this theory, we demonstrates EPC-induced giant PMMs leading to large phonon Zeeman splittings in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models.
+Building on these splittings, we propose the inertially decoupled model to search for intrinsic phonon Chern states in realistic magnetic materials.
+Furthermore, through first-principles calculations, we propose promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$.
+Our work establishes a ab initio theoretical foundation for investigating EPC-induced phonon magnetic properties and TRS-breaking topological phonon spectra in realistic materials, which could pave the way for futher developments of chiral phonon physics.
+Ab initio theory for phonon Zeeman energy.
 Previous theories regarding the magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals and are only effective for specific high-symmetry phonon wave vectors [7, 9, 10, 12].
-However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop a unified theory that can comprehensively describe phonon behaviors in insulators, semimetals, and metals.
+However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop an ab initio theory that can comprehensively describe phonon magnetic properties in insulators, semimetals, and metals.
+Instead of focusing on calculating effective PMM, we choose to directly calculate experimentally observable Zeeman splittings.
 We begin with a TRS-breaking perturbation phonon Hamiltonian, $\delta H=-\sum_{ab}\eta_{ab}p_{a}u_{b}$, known as the Raman spin-phonon interaction [43], where $\eta_{ab}$ is a real antisymmetric matrix that satisfies $\eta_{ab}=-\eta_{ba}$.
 Here, $u_{a}=\sqrt{m}x_{a}$ represents the renormalized atomic displacement coordinate and $p_{a}$ represents the canonical momentum related to $u_{a}$, where $m$ denotes the atomic mass.
-By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$.
-Here, we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$.
-Therefore, the Zeeman energy of the chiral phonon can be regarded as an effective linear response to its angular momentum.
-We notice that $\bm{\eta}$ can be deduced from the TRS-breaking self-energy correction of chiral phonons driven by EPC.
-Remarkably, within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$.
+By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$, where we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$.
+Based on the expression of phonon Zeeman energy, we propose that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system.
+Therefore, the Zeeman energy of chiral phonon can be regarded as an effective linear response to its angular momentum.
+Remarkably, $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons induced by EPC.
+Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$.
 The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig.
-1(b)(see the supplement material (SM)) [45],
+1(b)[see the supplement material (SM)] [44],
 $\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (1)
@@
 Consequently, we can conclude that EPC-driven large phonon Zeeman splittings are most likely to occur in narrow-gap semiconductors, semimetals and metals comprising light elements.
-Large phonon Zeeman splittings suggest that phonons exhibit significant phonon magnetic moments.
+Large phonon Zeeman splittings suggest that phonons exhibit significant PMM.
 To derive the PMM, we taking the derivative $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}$ and the phonon Zeeman energy is further rewritten as $E_{\text{Z}}=-\bm{B}\cdot\bm{\mu}^{\text{ph}}$, where $\bm{B}$ represents the effective magnetic field, and $\bm{\mu}^{\text{ph}}$ denotes the PMM (see SM  [44]),
@@
 Candidate materials.
-Building upon the above theoretical framework, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [46, 47] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons.
+Building upon the our theory, we developed a dedicated computational module within Quantum ESPRESSO [51, 52] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons.
 Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from dynamical matrix through near-degenerate perturbation theory (see SM [44]), yielding TRS-broken phonon dispersions.
@@
 Thus, the gap determines the maximum operating temperature for such a detector.
-Fig.
-4(b) lists the transfered energy ranges and gaps for all five candidate materials (see SM [45] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$.
-This energy range corresponds to light-mass dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates.
+Figure 4(b) lists gaps and transfered energy ranges for proposed candidate materials (see SM [44] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$.
+This energy range corresponds to low-energy dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates [63, 64].
 Figure 4: Detectors of neutral particles based on the intrinsic phonon Chern state.
@@
 (b) Transfered energy ranges and phonon Chern gaps for all the layered candidate materials.
-In summary, we have established a unified ab initio framework enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems.
-Under this framework, we propose an inertially decoupled model to design intrinsic phonon Chern states, and identify five candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz.
-Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through phonon edge current transduction.
-These advances simultaneously open broad pathways for revealing new phonon magnetic effects in quantum magnetic materials.
+In summary, we have established an ab initio theory enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems.
+Based on our theory, we build the inertially decoupled model to seek intrinsic phonon Chern states, and propose candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz.
+Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through robust phonon edge currents.
+These advances could open broad pathways for investigating phonon magnetic effects in quantum materials.
 Acknowledgements.

### v3

@@
-Ab Initio theory of Electron-phonon-coupling-induced Giant Magnetic Moments of Chiral Phonons in Magnetic Materials
+Ab Initio Theory of Large Phonon Magnetic Moments Induced by Electron-Phonon Coupling in Magnetic Materials
 Abstract
 Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention.
-However, a long-standing critical issue in this field is the lack of ab initio method to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC).
-Here, we resolve this challenge by developing an ab initio theory for calculating EPC-induced phonon magnetic properties, applicable to both insulating and metallic materials.
-Based on this theory, we demonstrate EPC-induced giant phonon magnetic moments and resulting phonon Zeeman splittings in magnetic metals, which are orders of magnitude larger than classical predictions from point-charge models.
-Interestingly, these splittings could open observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states.
-Through first-principles calculations, we first propose candidate materials with such intrinsic phonon Chern states hosting robust edge phonon currents which may be applied to detecting neutral particles such as dark matter particles.
-Our work not only establishes a theoretical foundation for EPC-induced phonon magnetic properties, but also enables the ab initio calculation of long-sought TRS-breaking phonon spectra throughout the Brillouin zone in realistic materials.
+However, a long-standing critical issue in this field is the lack of an approach to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC) in realistic materials.
+Here, based on the linear response framework, we develop an ab initio theory for calculating EPC-induced magnetic properties of phonons, applicable to both insulating and metallic materials.
+Then, we evidently demonstrate EPC-induced large phonon magnetic moments and phonon Zeeman splittings resulting from spin-phonon coupling in magnetic metals.
+Interestingly, these splittings can open substantial topologically nontrivial phonon gaps, generating intrinsic phonon Chern states.
+Further, by constructing an inertially decoupled lattice model, we predict candidate materials exhibiting such intrinsic phonon Chern states with robust edge phonon currents which are proposed to detect neutral particles, such as dark matter particles.
+Our work enables ab initio calculations for EPC-induced magnetic properties of phonons and long-sought magnetic phonon spectra.
 Introduction.
 Chiral phonons, collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and Niu [1, 2] and subsequently verified experimentally in monolayer WSe2 [3].
-Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and experiments [15, 16, 17, 18, 19, 20, 21, 22, 23].
-Of particular importance are phonon magnetic moments (PMMs) of chiral phonons and their associated phonon Zeeman splittings, playing a crucial role in chiral-phonon-involved effects [1, 24, 25, 26, 27].
-However, classical models based on point-charge approximations yield extremely small PMM values due to the significant large mass of nucleus  [9].
-While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [7, 9, 12], an ab initio method capable of accurately calculating EPC-induced PMMs in general materials is still lacking.
-Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-induced PMM.
-In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions, which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons.
-The classical Zeeman splitting of chiral phonons deduced from the point-charge model is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger.
+Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] and experiments [19, 20, 21, 22, 23, 24, 25, 26, 27].
+Of particular importance are phonon magnetic moments (PMMs) and phonon Zeeman splittings, which play a crucial role in chiral-phonon-involved effects, such as ultrafast Einstein-de Haas Effect [20] and Barnett effect [22, 28].
+However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below experimental measurements [11, 19].
+While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [8, 11, 15], an ab initio approach capable of accurately calculating EPC-induced PMMs in realistic materials is lacking.
+Figure 1: Schematic of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-enhanced PMM.
+In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions (black arrows), which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons.
+The classical phonon Zeeman splitting deduced from the point-charge model ($Z_{\text{eff}}$ is the effective charge) is negligible as a result of the significant atomic mass, whereas the EPC-induced phonon Zeeman splitting (b) is orders of magnitude larger.
 This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
 Besides magnetic moments, phonons can also exhibit Hall effects [29, 30, 31, 32] and topological characteristics, such as Berry curvature and Chern numbers [33, 34, 35].
-When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents with significantly enhanced detectability compared to conventional phonon Hall effects.
-Nevertheless, TRS-breaking external fields for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations [35, 36].
-Consequently, phonon Chern states driven by the phonon Zeeman splitting in magnets without external fields are highly desirable.
-However, predicting such intrinsic topological phonon states in magnets remains challenging due to the inability to perform ab initio calcualtions of TRS-broken phonon spectra [37, 38, 39, 40, 41, 42].
-To resolve these challenges, we establish an ab initio theory for determining EPC-induced phonon magnetic properties, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals.
-Our theory relies on treating the angular momentum of chiral phonon as an effective perturbative field to the electron-phonon coupled system, with the phonon Zeeman energy interpreted as the response to this perturbation.
-Based on this theory, we demonstrates EPC-induced giant PMMs leading to large phonon Zeeman splittings in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models.
-Building on these splittings, we propose the inertially decoupled model to search for intrinsic phonon Chern states in realistic magnetic materials.
-Furthermore, through first-principles calculations, we propose promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$.
-Our work establishes a ab initio theoretical foundation for investigating EPC-induced phonon magnetic properties and TRS-breaking topological phonon spectra in realistic materials, which could pave the way for futher developments of chiral phonon physics.
-Ab initio theory for phonon Zeeman energy.
-Previous theories regarding the magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals and are only effective for specific high-symmetry phonon wave vectors [7, 9, 10, 12].
-However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop an ab initio theory that can comprehensively describe phonon magnetic properties in insulators, semimetals, and metals.
-Instead of focusing on calculating effective PMM, we choose to directly calculate experimentally observable Zeeman splittings.
-We begin with a TRS-breaking perturbation phonon Hamiltonian, $\delta H=-\sum_{ab}\eta_{ab}p_{a}u_{b}$, known as the Raman spin-phonon interaction [43], where $\eta_{ab}$ is a real antisymmetric matrix that satisfies $\eta_{ab}=-\eta_{ba}$.
-Here, $u_{a}=\sqrt{m}x_{a}$ represents the renormalized atomic displacement coordinate and $p_{a}$ represents the canonical momentum related to $u_{a}$, where $m$ denotes the atomic mass.
-By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$, where we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$.
-Based on the expression of phonon Zeeman energy, we propose that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system.
-Therefore, the Zeeman energy of chiral phonon can be regarded as an effective linear response to its angular momentum.
-Remarkably, $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons induced by EPC.
-Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$.
+When the time-reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents.
+Nevertheless, TRS-breaking external fields for phonons, such as strong magnetic fields or Coriolis forces, often face experimental limitations [36, 37].
+Consequently, phonon Chern states driven by the phonon Zeeman splitting resulting from spin-phonon coupling (SPC) in magnets without requiring external fields are highly desirable.
+However, it is challenging to predict such intrinsic topological phonon states in magnets due to the inability to perform ab initio calculations of TRS-breaking phonon spectra [38, 39, 40, 41, 42, 43].
+To resolve these challenges, we establish an ab initio theory for determining EPC-induced magnetic properties of phonons, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals.
+Based on this theory, we demonstrate that EPC-induced PMMs can lead to large phonon Zeeman splittings through SPC in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models [Fig.
+1(a)].
+Considering that these splittings could open significant topological gaps in phonon dispersions [Fig.
+1(b)], we propose an inertially decoupled lattice model to realize intrinsic phonon Chern states in realistic materials.
+Furthermore, through first-principles calculations, we predict a family of promising candidate materials exhibiting intrinsic phonon Chern states.
+Our work establishes an ab initio theoretical foundation for investigating EPC-induced magnetic properties of phonons, which paves the way for further developments of phonon physics.
+Ab initio theory for phonon Zeeman energy and phonon magnetic moment.
+Previous theories regarding magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals, and are restricted to the $\Gamma$ point [8, 11, 13, 15].
+Since EPC is typically much more significant in metallic systems, we develop an ab initio theory that can generally calculate EPC-induced magnetic properties of phonons across the whole Brillouin zone for both metals and insulators, which is also compatible with previous theories [see the supplemental material (SM) [44]].
+We will demonstrate our theory as follows.
+Instead of directly calculating the magnetic moment of a single phonon, we focus on experimentally observable phonon Zeeman energy induced by the coupling between PMM and spins of electrons.
+We start from the total phonon Hamiltonian including the TRS-breaking perturbation as
+$H=\sum_{l,a}\frac{1}{2}p_{l,a}^{2}+\frac{1}{2}\sum_{l,ab}D_{l,ab}u_{l,a}u_{l,b}-\sum_{l,ab}\eta_{l,ab}p_{l,a}u_{l,b},$ (1)
+where $a,b=x,y,z$ represent the Cartesian coordinates, $l=1,2,...N$ denotes the lattice site for a system containing $N$ atoms, and $u_{l,a}=\sqrt{m}x_{l,a}$ represents the renormalized atomic displacement with $m$ being the atomic mass.
+The last term in Eq.
+(1), denoted as $\delta H$, describes the TRS-breaking perturbation resulting from SPC [1], and $\eta_{l,ab}$ is a $3\times 3$ real antisymmetric matrix satisfying $\eta_{l,ab}=-\eta_{l,ba}$.
+In addition, $p_{l,a}=\dot{u}_{l,a}+\sum_{l,b}\eta_{l,ab}u_{l,b}$ represents the canonical momentum [44].
+By identifying the phonon angular momentum as $s^{\text{ph}}_{l,c}=\sum_{ab}\epsilon_{abc}u_{l,a}p_{l,b}$ (c is the direction of angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express $\delta H$ as phonon Zeeman energy in a compact form $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}=\sum_{l,ab}\eta_{l,c}s^{\text{ph}}_{l,c}$, where we have defined $\eta_{l,c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{l,ab}$.
+Notably, we find that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system, and thus the Zeeman energy of chiral phonon can be effectively regarded as a linear response.
+Importantly, we will show that $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons.
+Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where $\chi_{c}^{-}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$ [see SM [44]], indicating that the spin polarization of electrons leads to Zeeman splittings of chiral phonons.
 The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig.
-1(b)[see the supplement material (SM)] [44],
-$\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (1)
-Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively.
-$N$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$.
+1(b),
+$\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N_{\bm{k}}}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (2)
+Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively, and $\omega$ is the phonon frequency.
+$N_{\bm{k}}$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$.
 $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [45, 46].
-Notably, all parameters in Eq.
-(1) can be directly derived from first-principles calculations, obviating the need for empirical inputs or additional assumptions.
-As a result, the previously inaccessible TRS-breaking phonon spectra driven by EPC in both insulating and metallic systems is now accessible across the whole Brillouin zone.
-Phonon Zeeman splitting and enhanced phonon $g$-factor.
-We propose the essential conditions for realizing significant phonon Zeeman splittings by EPC.
-Since the primary determinant of the phonon Zeeman splitting is the susceptibility $\bm{\chi}^{-}$, there are two key points according to Eq.
+It should be emphasized that all parameters in Eq.
+(2) can be directly obtained from first-principles calculations, without empirical inputs or additional assumptions.
+A non-zero $\bm{\chi}^{-}$ requires spin-polarized electronic states, for example, in ferromagnets.
+To realize large $\bm{\chi}^{-}$, there are two key points according to Eq.
 (2): (i) The contribution of EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is substantial, and materials composed of light atoms tend to possess strong $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$.
-(ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, the parametric sensitivity in Eq.
-(1) leads to an enhancement of $\bm{\chi}^{-}$ [44].
-Consequently, we can conclude that EPC-driven large phonon Zeeman splittings are most likely to occur in narrow-gap semiconductors, semimetals and metals comprising light elements.
-Large phonon Zeeman splittings suggest that phonons exhibit significant PMM.
-To derive the PMM, we taking the derivative $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}$ and the phonon Zeeman energy is further rewritten as $E_{\text{Z}}=-\bm{B}\cdot\bm{\mu}^{\text{ph}}$, where $\bm{B}$ represents the effective magnetic field, and $\bm{\mu}^{\text{ph}}$ denotes the PMM (see SM  [44]),
-$\mu^{\text{ph}}_{c}=-g^{\text{ph}}_{c}\frac{\mu_{B}}{\hbar}s^{\text{ph}}_{c}.$ (2)
-Here, $\mu_{B}$ denotes the Bohr magneton, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon, where $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$, $g^{e}_{c}$ is the $g$-factor of electrons along $c$ direction involved in EPC, and the dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor.
-This amplification factor is primarily influenced by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electronic states, and it can significantly exceed unity.
-This enhancement allows the $g$-factor of the chiral phonon to the order of $g^{e}_{c}$.
-This EPC-enhanced PMM can be intuitively understood through the following mechanism.
-As chiral phonons drive atomic cyclotron motion, a portion of its angular momentum is transferred to the electrons via EPC.
-Owing to the considerably small mass of electrons, the angular momentum transfer gives rise to strong loop electron currents, thereby resulting in a significant PMM.
-In contrast, within the classical framework, the phonon’s $g$-factor is constrained to the order of $10^{-5}\sim 10^{-4}g^{e}_{c}$ due to heavy ionic masses compared to the electron mass [9, 47].
+(ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, it leads to an enhancement of $\bm{\chi}^{-}$.
+Consequently, we can conclude that EPC-induced large phonon Zeeman splittings are most likely to occur in ferromagnetic narrow-gap semiconductors, semimetals and metals containing light elements.
+Then, PMM can be derived from the above phonon Zeeman energy as (see SM [44])
+$\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}=-g^{\text{ph}}_{c}\frac{\mu_{B}}{\hbar}s^{\text{ph}}_{c}.$ (3)
+Here, $\mu_{B}$ denotes the Bohr magneton of electron, $B_{c}$ is the effective magnetic field, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon (see SM [44]), where $E_{F}$ is the Fermi energy and $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$.
+$g^{e}_{c}$ is the $g$-factor along $c$ direction of electrons involved in EPC.
+Equation (3) enables first-principles calculations of PMM.
+The dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor dominated by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electron states near the Fermi energy [Eq.
+(S22) in SM [44]], and it allows the $g$-factor of the chiral phonon to reach the order of $g^{e}_{c}$, leading to a large PMM.
+This enhancement of PMM originates from electron loop currents induced by the EPC-mediated angular momentum transfer from ions to electrons, where electron loop currents are significantly enhanced owing to the tiny electron mass.
+In contrast, within the classical framework, the phonon’s $g$-factor is constrained to $10^{-5}\sim 10^{-4}$ order of $g^{e}_{c}$ due to large ionic masses compared to the electron mass [11, 47].
 Figure 2: Inertially decoupled model (IDM) and intrinsic phonon Chern states.
-(a) Schematic of optical chiral phonon modes: When $m\sim M$, both atoms contribute to the high-frequency model (up); in the IDM condition ($m\ll M$), the optical model is primarily contributed by light atoms (down).
-(b) Phonon dispersions of nonmagnetic honeycomb lattice and magnetic triangle lattice.
-(c-e) TRS-breaking phonon dispersions of 2D/3D IDM structures: the 2D IDM by combining the honeycomb and triangular lattices (c,d), and the 3D IDM by stacking 2D IDM along the $z$-direction (e).
-The color gradient from violet and yellow represents the contributions from the honeycomb and triangular lattice, and the colored gradient from blue and red represents the phonon angular momentum magnitude.
-(f) Edge states (cyan/magenta) of IDM under open boundaries.
+(a) Schematic of chiral phonon modes in the high frequency regime: When $m\sim M$, both atoms contribute to high-frequency modes (upper panel); in the IDM condition ($m\ll M$), high-frequency modes are primarily contributed by light atoms (lower panel).
+(b) Phonon dispersions of nonmagnetic honeycomb lattice and ferromagnetic triangular lattice.
+(c-e) TRS-breaking phonon dispersions of the 2D IDM by combining the honeycomb and triangular lattices (c,d) and the 3D IDM by stacking 2D IDM along the $z$-direction (e).
+(f) Edge states (cyan/magenta) of the 2D IDM.
 Parameters are detailed in the SM [44].
 Intrinsic phonon Chern states and the inertially decoupled model.
-The large phonon Zeeman splitting can pave the way toward a phonon Chern state through the phonon band inversion [44].
+The accurate calculation of phonon Zeeman splitting enables the determination of magnetic phonon spectra throughout the Brillouin zone, which is indispensable for exploring magnetic phonon topology.
 Following the analogy to the Chern state in the Haldane model [48], we adopt the two-dimensional (2D) honeycomb lattice [49, 50, 51] as an illustrative example.
-When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $m_{I}$ and $m_{T}$, respectively.
-It has been shown that when $|m_{T}|>|m_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [43].
-However, realizing such phonon Chern states in realistic materials hosting honeycomb lattices remains challenging.
-Although Coriolis forces and strong magnetic fields can produce a nonzero $m_{T}$ [43], these approaches either require large external fields or complex experimental measurements.
-In contrast, the EPC-enhanced phonon Zeeman splitting could offer a sizable $m_{T}$ in magnets without the need of external fields.
-To further satisfy the condition $|m_{T}|>|m_{I}|$, it is necessary to minimize $|m_{I}|$.
-Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|m_{I}|=0$ would be the ideal candidate.
+When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $\Delta_{I}$ and $\Delta_{T}$, respectively.
+It has been shown that when $|\Delta_{T}|>|\Delta_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [49].
+However, it is challenging to realize such phonon Chern states in realistic materials with honeycomb lattices.
+Although Coriolis forces and strong magnetic fields can produce a nonzero $\Delta_{T}$ [49], these approaches either require large external fields or complex experimental measurements.
+In contrast, the phonon Zeeman splitting could offer a sizable $\Delta_{T}$ in magnets without the need of external fields.
+To further satisfy the condition of $|\Delta_{T}|>|\Delta_{I}|$, it is necessary to minimize $|\Delta_{I}|$.
+Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|\Delta_{I}|=0$ would be the ideal candidate.
 Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered.
-Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $m_{I}$, which will make the condition $|m_{T}|>|m_{I}|$ difficult to satisfy.
+Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $\Delta_{I}$, which will make the condition $|\Delta_{T}|>|\Delta_{I}|$ difficult to satisfy.
 To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference.
@@
 2(b,c).
-Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions mainly originate from the light and heavy atoms, respectively, as shown in Figs.
+Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions are mainly contributed by the light and heavy atoms, respectively, as shown in Figs.
 2(a,c), which is termed “inertially decoupled” (see SM [44] for details).
-Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $m_{I}$ mass at both the $K$ and $K^{\prime}$ valleys.
-Meanwhile, the magnetic $M$ atoms breaks TRS and provide effective magnetic fields to generate large Zeeman splittings of chiral phonons via EPC, producing the desired large $m_{T}$ mass.
-The large $m_{T}$ opens a topologically nontrivial phonon gap (seen in Fig.
-2(c,d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by the topological edge states crossing the phonon gap, as shown in Fig.
+Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $\Delta_{I}$ mass at both the $K$ and $K^{\prime}$ valleys.
+Meanwhile, the magnetic $M$ atoms break TRS to generate large Zeeman splittings of chiral phonons via SPC, producing the desired large $\Delta_{T}$ mass.
+The large $\Delta_{T}$ opens a topologically nontrivial phonon gap (seen in Fig.
+2(d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by topologically protected edge states crossing the phonon gap, as shown in Fig.
 2(f).
 The inertially decoupled model is also applicable to three-dimensional (3D) magnets.
-Interestingly, when the nonmagnetic honeycomb lattice (the $m$ sublayer) and the magnetic triangular lattice (the $M$ sublayer) are stacked alternately along the $z$-direction, each $M$ sublayer effectively acts as a buffer layer that nearly decouples the phonon dispersions of adjacent $m$ sublayers in the high-frequency regime.
+Interestingly, when the nonmagnetic honeycomb lattice (the $m$ sublayer) and the ferromagnetic triangular lattice (the $M$ sublayer) are stacked alternately along the $z$-direction, each $M$ sublayer effectively acts as a buffer layer that nearly decouples the phonon dispersions of adjacent $m$ sublayers in the high-frequency regime.
 A representative phonon dispersion of such a 3D bulk model is shown in Fig.
 2(e).
-Notably, since each $m$ sublayer realizes a Chern state with $|C|=1$, the system exhibits an intrinsic 3D phonon Chern state.
-This configuration hosts multiple chiral phonon edge states resulting from a total Chern number $|C|=N_{z}$, where $N_{z}$ is the number of $m$ sublayers.
+Notably, since each $m$ sublayer realizes a Chern state with $|C|=1$, the system exhibits a 3D phonon Chern state.
+This configuration hosts multiple unidirectional phonon edge states resulting from a total Chern number $|C|=N_{z}$, where $N_{z}$ is the number of $m$ sublayers.
 Figure 3: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$.
-The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$.
-(c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the TRS breaking, respectively.
-(e) The phonon dispersion of $\mathrm{EuSi_{2}}$ under open-boundary conditions.
-(f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the TRS breaking, respectively.
+(a,b) The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$.
+(c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the SPC, respectively.
+(e) The phonon dispersion of bulk $\mathrm{EuSi_{2}}$ under open-boundary conditions.
+(f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the SPC, respectively.
 (h) The phonon dispersion of the sandwiched slab under open-boundary conditions.
-The color gradient from red to blue represents the magnitude of the phonon angular momentum.
-Cyan and magenta label the surface or edge modes originating from opposite boundaries, respectively.
+The color gradient from red to blue represents the phonon angular momentum.
+Cyan and magenta colors in (c) and (h) label topological boundary modes residing on opposite boundaries.
 Candidate materials.
-Building upon the our theory, we developed a dedicated computational module within Quantum ESPRESSO [51, 52] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons.
-Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from dynamical matrix through near-degenerate perturbation theory (see SM [44]), yielding TRS-broken phonon dispersions.
-Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal $P6/mmm$ magnets, specifically $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [57], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [58], and $\mathrm{CrB_{2}}$ [59, 60], exhibiting intrinsic phonon Chern states via EPC-enhanced Zeeman splittings ($\Delta_{\text{z}}=2|E_{\text{Z}}|$).
-As summarized in Table I, they all have large observable phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\gamma$ is phonon linewidth resulting from the EPC [61, 62].
+Building upon our theory, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density-functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons.
+Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from the dynamical matrix through near-degenerate perturbation theory (see SM [44] for details), yielding TRS-breaking phonon dispersions.
+Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal ($P6/mmm$) magnets exhibiting intrinsic phonon Chern states, namely, $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [59], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [60], and $\mathrm{CrB_{2}}$ [61, 62].
+Remarkably, they all have large observable topological phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\Delta_{\text{z}}=2|E_{\text{Z}}|$ is the phonon Zeeman splitting and $\gamma$ is the phonon linewidth resulting from EPC [63, 64], as summarized in Table I.
 Note that the Zeeman splitting should be larger than the phonon linewidth to produce an observable phonon gap.
@@
 The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs.
-3(a, e).
-Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a magnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction.
+3(a, b).
+Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a ferromagnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction.
 The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately $28\,u$ and $152\,u$ ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$.
@@
 Our first-principles calculations reveal that the EPC induces a pronounced Zeeman splitting of chiral phonons in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs.
-3(c, d)], opening a sizable full gap ($\approx 0.35$ THz) throughout the Brillouin zone.
+3(c, d)], opening a sizable full topological gap ($\approx 0.35$ THz) throughout the Brillouin zone.
 Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state.
@@
 3(e).
-To achieve a 2D Chern state, we construct a sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig.
+To achieve a 2D Chern state, we construct a dynamically-stable sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig.
 3(b).
 By comparing Figs.
-2(f) and 2(g), we can see that the Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ in the phonon dispersion of this sandwiched slab, and the accompanying single pair of edge modes traversing the phonon gap, as shown in Fig.
-3(h), provide definitive evidence for intrinsic phonon Chern states.
-Table 1: The direct phonon gaps at K points for the five candidate materials, where $\gamma$ and $\Delta_{\text{z}}$ represent the phonon linewidth and the phonon Zeeman splitting, respectively.
+2(f) and 2(g), we can see that the phonon Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ featuring a pair of edge modes traversing the phonon gap, as shown in Fig.
+3(h).
+This provides definitive evidence for intrinsic phonon Chern states.
+Table 1: The direct phonon gaps at K points for the candidate materials, where $\gamma$ and $\Delta_{\text{z}}$ represent the phonon linewidth and the phonon Zeeman splitting, respectively.
 The observable phonon gap $\Delta_{\text{ph}}$ is defined as $\Delta_{\text{z}}-\gamma$.
@@
 Discussion and conclusion.
-The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring extreme external strong magnetic fields, Coriolis forces, or stress gradients.
-One potential advantage may be its ability to detect neutral particles by converting them into phonon edge currents.
+The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring strong external magnetic fields, Coriolis forces, or stress gradients [2, 49].
+One potential advantage may be its ability to detect neutral particles by converting them into topological phonon edge currents.
 Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging.
-However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within that of the first “conduction” band above the topological gap in the phonon dispersion, this process may potentially excite the corresponding phonon modes [63, 64], driving the system out of equilibrium.
-As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via edge states, producing a edge phonon current immune to scattering[Fig.
+However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within the first “conduction” band above the topological gap in the phonon dispersion, corresponding phonon modes can be excited [65, 66], driving the system out of equilibrium.
+As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via topologically protected edge states, manifesting as a unidirectional edge phonon current immune to scattering [Fig.
 4(a)].
-This robust edge current is significantly more measurable than conventional phonon excitations dispersed throughout the material.
+This robust edge current is significantly more measurable than conventional delocalized bulk phonon excitations in materials.
 However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents.
-Thus, the gap determines the maximum operating temperature for such a detector.
-Figure 4(b) lists gaps and transfered energy ranges for proposed candidate materials (see SM [44] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$.
-This energy range corresponds to low-energy dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates [63, 64].
-Figure 4: Detectors of neutral particles based on the intrinsic phonon Chern state.
+Thus, a large phonon Chern gap is highly desired to elevate the working temperature.
+Figure 4(b) lists phonon Chern gaps and transferred energy ranges of proposed candidate materials (see SM [44] for details), with transferred energy spanning approximately $20\sim 110\,\mathrm{meV}$.
+This energy range corresponds to low-energy dark matter which recently attracts much attention, therefore our work provides a potential detection mechanism for this category of dark matter candidates [65, 66].
+Figure 4: Detection of neutral particles based on the intrinsic phonon Chern state.
 (a) Schematic illustrations of the detection process.
-Neutral particles excite phonon modes in high-frequency “conduction” bands.
-These modes then gradually relax to lower-frequency regimes via chiral edge phonon channels.
-(b) Transfered energy ranges and phonon Chern gaps for all the layered candidate materials.
-In summary, we have established an ab initio theory enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems.
-Based on our theory, we build the inertially decoupled model to seek intrinsic phonon Chern states, and propose candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz.
-Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through robust phonon edge currents.
+Left: Neutral particles (red beam) excite phonon modes (yellow circles in the lower panel) in high-frequency bands.
+Right: These phonon modes gradually relax to lower-frequency regimes via topologically protected edge phonon channels (orange and green lines across the gap), generating robust unidirectional phonon edge currents (orange and green arrows).
+(b) Transferred energy ranges and phonon Chern gaps from proposed candidate materials [Tab.
+S2 in SM [44]].
+In summary, we have established an ab initio theory enabling systematic predictions of EPC-induced magnetic properties of phonons and TRS-breaking phonon dispersions for insulators, semimetals and metals.
+Based on our theory, we construct the inertially decoupled model to seek intrinsic phonon Chern states in realistic materials, and propose candidates exhibiting EPC-induced large topological phonon gaps exceeding 0.3 THz.
+Furthermore, we demonstrate the potential capability of these topological phonon states to detect neutral particles through robust phonon edge currents.
 These advances could open broad pathways for investigating phonon magnetic effects in quantum materials.

### v4

@@
-Ab Initio Theory of Large Phonon Magnetic Moments Induced by Electron-Phonon Coupling in Magnetic Materials
+Ab Initio Theory of Phonon Magnetic Moment Induced by Electron-Phonon Coupling in Magnetic Materials
 Abstract
-Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention.
+Circularly polarized phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention.
 However, a long-standing critical issue in this field is the lack of an approach to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC) in realistic materials.
 Here, based on the linear response framework, we develop an ab initio theory for calculating EPC-induced magnetic properties of phonons, applicable to both insulating and metallic materials.
-Then, we evidently demonstrate EPC-induced large phonon magnetic moments and phonon Zeeman splittings resulting from spin-phonon coupling in magnetic metals.
-Interestingly, these splittings can open substantial topologically nontrivial phonon gaps, generating intrinsic phonon Chern states.
-Further, by constructing an inertially decoupled lattice model, we predict candidate materials exhibiting such intrinsic phonon Chern states with robust edge phonon currents which are proposed to detect neutral particles, such as dark matter particles.
-Our work enables ab initio calculations for EPC-induced magnetic properties of phonons and long-sought magnetic phonon spectra.
+Our method can precisely calculate phonon Zeeman splittings in magnetic metals with significant EPC, as demonstrated by the remarkable agreement with recent experimental observations of phonon Zeeman splitting in the ferromagnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$.
+In addition, the long-sought magnetic phonon spectra across the entire Brillouin zone are obtained, facilitating the study of magnetic phonon transport and topology.
+Specifically, by constructing an inertially decoupled lattice model, we propose candidate materials exhibiting intrinsic phonon Chern states with robust unidirectional edge phonon currents.
+Our work paves the way for investigating novel phonon phenomena in magnetic quantum materials.
 Introduction.
-Chiral phonons, collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and Niu [1, 2] and subsequently verified experimentally in monolayer WSe2 [3].
-Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] and experiments [19, 20, 21, 22, 23, 24, 25, 26, 27].
-Of particular importance are phonon magnetic moments (PMMs) and phonon Zeeman splittings, which play a crucial role in chiral-phonon-involved effects, such as ultrafast Einstein-de Haas Effect [20] and Barnett effect [22, 28].
+Circularly polarized phonons, collective vibrations stemming from the circular motion of atoms in solids that carry nonzero angular momentum [1, 2], are referred to as axial phonons [3].
+Recently, there has been a surge of research interest in circularly polarized phonons [4, 5], covering both theoretical investigations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] and experiments [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
+Of particular importance is the phonon magnetic moment (PMM), which is derived from phonon energy shift in a magnetic field [14].
+PMM plays a crucial role in various effects involving circularly polarized phonons, such as ultrafast Einstein-de Haas Effect [30] and Barnett effect [32, 40].
 However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below experimental measurements [16, 29].
-While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [8, 11, 15], an ab initio approach capable of accurately calculating EPC-induced PMMs in realistic materials is lacking.
+While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [14, 16], to our knowledge, there is no comprehensive ab initio framework to quantitatively predict EPC-induced PMMs in realistic materials.
+Previous theories describing EPC-induced PMM are only applicable to either insulators or semimetals [14, 15, 16, 13, 19, 21, 22, 20], and more importantly, most of them rely on empirical parameters and, consequently, deviate from ab initio approach.
 Figure 1: Schematic of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-enhanced PMM.
-In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions (black arrows), which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons.
+In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions (black arrows), which can be recombined into left-circularly (red) and right-circularly (blue) polarized phonons.
 The classical phonon Zeeman splitting deduced from the point-charge model ($Z_{\text{eff}}$ is the effective charge) is negligible as a result of the significant atomic mass, whereas the EPC-induced phonon Zeeman splitting (b) is orders of magnitude larger.
-This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
-Besides magnetic moments, phonons can also exhibit Hall effects [29, 30, 31, 32] and topological characteristics, such as Berry curvature and Chern numbers [33, 34, 35].
-When the time-reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents.
-Nevertheless, TRS-breaking external fields for phonons, such as strong magnetic fields or Coriolis forces, often face experimental limitations [36, 37].
-Consequently, phonon Chern states driven by the phonon Zeeman splitting resulting from spin-phonon coupling (SPC) in magnets without requiring external fields are highly desirable.
-However, it is challenging to predict such intrinsic topological phonon states in magnets due to the inability to perform ab initio calculations of TRS-breaking phonon spectra [38, 39, 40, 41, 42, 43].
-To resolve these challenges, we establish an ab initio theory for determining EPC-induced magnetic properties of phonons, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals.
-Based on this theory, we demonstrate that EPC-induced PMMs can lead to large phonon Zeeman splittings through SPC in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models [Fig.
+Meanwhile, the coupling between PMM and magnetic order through spin-phonon interactions yields rich phenomena, such as phonon Zeeman splittings, phonon Hall effects [41, 42, 43, 44], and magnetic topological phonon states [45, 5].
+For example, phonon Zeeman splitting induced by magnetic order was recently observed experimentally in the ferromagnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$ [46], yet a theory capable of quantitatively characterizing such Zeeman splittings is lacking.
+Moreover, investigations of phonon transport and phonon topology in magnetic systems usually require the full notion of the magnetic phonon spectra across the entire Brillouin zone.
+A prominent example is the phonon Chern state (phonon analogue of quantum anomalous Hall effect), which necessitates integrating Berry curvatures throughout the Brillouin zone to correctly obtain the topological Chern number [47, 48, 49, 50, 51].
+Although nonmagnetic phonon topology has been comprehensively investigated [52, 53, 54, 55, 56, 57], magnetic phonon topology remains largely unexplored due to challenges of calculating magnetic phonon spectra.
+To address these challenges, we establish an ab initio theory for determining EPC-induced magnetic properties of phonons, which enables ab initio calculations of previously inaccessible magnetic phonon spectra across material classes from insulators to metals.
+Based on this theory, we demonstrate that EPC-induced PMMs can lead to large phonon Zeeman splittings in magnetic metals [Fig.
+1(b)], which are orders of magnitude stronger than classical predictions from point-charge models [Fig.
 1(a)].
-Considering that these splittings could open significant topological gaps in phonon dispersions [Fig.
-1(b)], we propose an inertially decoupled lattice model to realize intrinsic phonon Chern states in realistic materials.
-Furthermore, through first-principles calculations, we predict a family of promising candidate materials exhibiting intrinsic phonon Chern states.
+Notably, our theory is further validated by the remarkable agreement with the experimental result of the phonon Zeeman splitting of $\mathrm{Co_{3}Sn_{2}S_{2}}$.
+By applying our theory, we predict a family of promising candidate materials with phonon Chern states according to the proposed inertially decoupled lattice model.
 Our work establishes an ab initio theoretical foundation for investigating EPC-induced magnetic properties of phonons, which paves the way for further developments of phonon physics.
 Ab initio theory for phonon Zeeman energy and phonon magnetic moment.
-Previous theories regarding magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals, and are restricted to the $\Gamma$ point [8, 11, 13, 15].
+Previous theories addressing magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals, and are restricted to the $\Gamma$ point [14, 16, 13, 19].
 Since EPC is typically much more significant in metallic systems, we develop an ab initio theory that can generally calculate EPC-induced magnetic properties of phonons across the whole Brillouin zone for both metals and insulators, which is also compatible with previous theories [see the supplemental material (SM) [58]].
@@
 Instead of directly calculating the magnetic moment of a single phonon, we focus on experimentally observable phonon Zeeman energy induced by the coupling between PMM and spins of electrons.
-We start from the total phonon Hamiltonian including the TRS-breaking perturbation as
+We start from the total phonon Hamiltonian including the time-reversal symmetry (TRS)-breaking perturbation as
 $H=\sum_{l,a}\frac{1}{2}p_{l,a}^{2}+\frac{1}{2}\sum_{l,ab}D_{l,ab}u_{l,a}u_{l,b}-\sum_{l,ab}\eta_{l,ab}p_{l,a}u_{l,b},$ (1)
-where $a,b=x,y,z$ represent the Cartesian coordinates, $l=1,2,...N$ denotes the lattice site for a system containing $N$ atoms, and $u_{l,a}=\sqrt{m}x_{l,a}$ represents the renormalized atomic displacement with $m$ being the atomic mass.
+where $a,b=x,y,z$ represent the Cartesian coordinates, l=1,2,Nl=1,2,...N denotes the lattice site for a system containing $N$ atoms, and $u_{l,a}=\sqrt{m}x_{l,a}$ represents the renormalized atomic displacement with $m$ being the atomic mass.
 The last term in Eq.
-(1), denoted as $\delta H$, describes the TRS-breaking perturbation resulting from SPC [1], and $\eta_{l,ab}$ is a $3\times 3$ real antisymmetric matrix satisfying $\eta_{l,ab}=-\eta_{l,ba}$.
+(1), denoted as $\delta H$, describes the TRS-breaking perturbation resulting from spin-phonon interaction [1], and $\eta_{l,ab}$ is a $3\times 3$ real antisymmetric matrix satisfying $\eta_{l,ab}=-\eta_{l,ba}$.
 In addition, $p_{l,a}=\dot{u}_{l,a}+\sum_{l,b}\eta_{l,ab}u_{l,b}$ represents the canonical momentum [58].
 By identifying the phonon angular momentum as $s^{\text{ph}}_{l,c}=\sum_{ab}\epsilon_{abc}u_{l,a}p_{l,b}$ (c is the direction of angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express $\delta H$ as phonon Zeeman energy in a compact form $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}=\sum_{l,ab}\eta_{l,c}s^{\text{ph}}_{l,c}$, where we have defined $\eta_{l,c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{l,ab}$.
-Notably, we find that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system, and thus the Zeeman energy of chiral phonon can be effectively regarded as a linear response.
-Importantly, we will show that $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons.
-Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where $\chi_{c}^{-}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$ [see SM [44]], indicating that the spin polarization of electrons leads to Zeeman splittings of chiral phonons.
-The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig.
-1(b),
+Notably, we find that phonon angular momentum can be treated as an effective perturbative field to the electron-phonon coupled system, and thus phonon Zeeman energy can be effectively regarded as a linear response.
+Importantly, we will show that $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of phonons. According to $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$, we have
+$\eta_{l,c}=\frac{\partial H}{\partial s^{\text{ph}}_{l,c}}=\frac{1}{p_{l,b}}\frac{\partial H}{\partial u_{l,a}}$ (2)
+where we have applied the chain derivative rule. Applying Kubo formula, we can get
+$\left<\eta_{l,c}\right>=-\frac{i}{\hbar}\int_{-\infty}^{t}dt^{\prime}u_{l,a}p_{l,b}\left\langle\left[\frac{-1}{p_{l,a}}\frac{\partial H}{\partial u_{l,b}}(t),\frac{1}{p_{l,b}}\frac{\partial H}{\partial u_{l,a}}(t^{\prime})\right]\right\rangle_{0}.$ (3)
+By taking Fourier transform of Eq.
+(3) and considering $p_{l,a}=-i\omega u_{l,a}$, we can obtain $\eta^{c}(\bm{q},\omega)$ ($\bm{q}$ is the phonon wave vector, and $\omega$ denotes phonon frequency) as
+$\eta_{c}(\bm{q},\omega)=\frac{1}{\hbar\omega}\int dte^{-i\omega t}\theta(t)\left\langle\left[\hat{V}_{a}(\bm{q},\omega),\hat{V}_{b}(-\bm{q},\omega)\right]\right\rangle_{0}.$ (4)
+Here, $\theta(t)$ is the Heaviside step function, and the EPC interaction operator $\hat{V}_{a}(\bm{q},\omega)=\sum_{l_{c}}z_{l_{c}}(\bm{q},\omega)\frac{\partial H}{\partial u_{l,a}}$, where $z_{l_{c}}(\bm{q},\omega)$ is the normalized contribution factor of the $l_{c}$-th atom for the phonon mode $(\bm{q},\omega)$[58].
+Now, we take electron’s spin degree of freedom into consideration.
+Taking the spin-quantization axis of electron as the c direction, we could get the relation that $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, with $\chi_{c}^{-}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$ (see SM [58] for details), indicating that the spin polarization of electrons leads to phonon Zeeman splittings.
+The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained by performing the Matsubara sum,
 $\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N_{\bm{k}}}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (5)
-Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively, and $\omega$ is the phonon frequency.
-$N_{\bm{k}}$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$.
+Here, $\bm{k}$ represents the wave vector of the electron, $N_{\bm{k}}$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic energy with band indices $\alpha$.
 $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [59, 60].
 It should be emphasized that all parameters in Eq.
-(2) can be directly obtained from first-principles calculations, without empirical inputs or additional assumptions.
+(5) can be directly obtained from ab initio calculations, without empirical inputs or additional assumptions.
 A non-zero $\bm{\chi}^{-}$ requires spin-polarized electronic states, for example, in ferromagnets.
@@
 (ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, it leads to an enhancement of $\bm{\chi}^{-}$.
-Consequently, we can conclude that EPC-induced large phonon Zeeman splittings are most likely to occur in ferromagnetic narrow-gap semiconductors, semimetals and metals containing light elements.
+Consequently, we can conclude that EPC-induced large phonon Zeeman splittings are most likely to occur in ferromagnetic narrow-gap semiconductors, semimetals, and metals containing light elements.
 Then, PMM can be derived from the above phonon Zeeman energy as (see SM [58])
 $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}=-g^{\text{ph}}_{c}\frac{\mu_{B}}{\hbar}s^{\text{ph}}_{c}.$ (6)
-Here, $\mu_{B}$ denotes the Bohr magneton of electron, $B_{c}$ is the effective magnetic field, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon (see SM [44]), where $E_{F}$ is the Fermi energy and $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$.
-$g^{e}_{c}$ is the $g$-factor along $c$ direction of electrons involved in EPC.
-Equation (3) enables first-principles calculations of PMM.
-The dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor dominated by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electron states near the Fermi energy [Eq.
-(S22) in SM [44]], and it allows the $g$-factor of the chiral phonon to reach the order of $g^{e}_{c}$, leading to a large PMM.
+Here, $\mu_{B}$ denotes the Bohr magneton of electron, $B_{c}$ is the effective magnetic field, and $g^{\text{ph}}_{c}=g^{e}\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E}|_{E=E_{F}}$ represents the effective phonon $g$-factor (see SM [58]), where $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$ with $g^{e}$ being the electronic $g$-factor involved in EPC.
+Equation (6) enables ab initio calculations of PMM.
+The dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E}|_{E=E_{F}}$ acts as an amplification factor dominated by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electron states near the Fermi energy [Eq.
+(S28) in SM [58]], and it allows the $g$-factor of the circularly polarized phonon to reach the order of $g^{e}$, leading to a large PMM.
 This enhancement of PMM originates from electron loop currents induced by the EPC-mediated angular momentum transfer from ions to electrons, where electron loop currents are significantly enhanced owing to the tiny electron mass.
-In contrast, within the classical framework, the phonon’s $g$-factor is constrained to $10^{-5}\sim 10^{-4}$ order of $g^{e}_{c}$ due to large ionic masses compared to the electron mass [11, 47].
+In contrast, within the classical framework, the phonon’s $g$-factor is constrained to $10^{-5}\sim 10^{-4}$ order of $g^{e}$ due to large ionic masses compared to the electron mass [16, 61].
+Based on the above theory, we develop a dedicated computational module within Quantum ESPRESSO [62, 63] incorporating density-functional perturbation theory with Hubbard corrections (DFPT+U) [59, 60] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons.
+Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from the dynamical matrix through near-degenerate perturbation theory (see SM [58] for details), yielding TRS-breaking phonon dispersions.
+Comparison with experiment.
+A recent experimental work reported the first observation of magnetic order induced circularly polarized phonons and their spontaneous splitting in the ferromagnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$ [46].
+As shown in Fig.
+2(a), $\mathrm{Co_{3}Sn_{2}S_{2}}$ belongs to centrosymmetric space group $\mathrm{R\bar{3}m}$, and exhibits out-of-plane ferromagnetic ordering below the Curie temperature ($\sim$175 K).
+Using helicity-resolved magneto-Raman scattering spectroscopy, a discernible splitting of the $E_{g}$ phonon modes at the $\Gamma$ point was observed in the ferromagnetic state, reaching up to $1.27~\mathrm{cm}^{-1}$($\sim$0.038 THz) at low temperatures (2 K).
+The two split modes correspond to phonon modes with opposite angular momenta, describing the right-handed and left-handed in-plane circular motion of S atoms, respectively [see Fig.
+2(c) insets].
+Although Ref.
+[46] provides an implicit EPC-based explanation of such a notable phonon Zeeman splitting, a direct and more convincing quantitative description is absent due to the lack of ab initio methods for magnetic phonon spectra.
+Now we show that our ab initio method could well reproduce the experimental result quantitatively.
+Giving full consideration to EPC effects, for the first time, we obtain the complete magnetic phonon dispersions of $\mathrm{Co_{3}Sn_{2}S_{2}}$ throughout the Brillouin zone (see the SM for more details [58]).
+Figure  2(b) presents the experimentally relevant magnetic phonon dispersion along the high-symmetry path $\Gamma$-T.
+Remarkably, our calculations clearly demonstrate a splitting of 0.042 THz ($\mu_{\text{ph}}=2.8\times 10^{-3}\mu_{B}$) at the $\Gamma$ point for the $E_{g}$ mode [Fig.
+2(c)], in good agreement with the observed value of 0.038 THz.
+Beyond this, we surprisingly find that the two circularly polarized phonon modes cross each other on the $\Gamma$-T path, forming a phonon Weyl point [Fig.
+2(c)].
+This magnetic phonon Weyl point [71] is protected by the $C_{3v}$ point group symmetry along $\Gamma$-T, which is previously inaccessible without obtaining the full magnetic phonon dispersions.
+Figure 2: Lattice structure and calculated magnetic phonon dispersions of $\mathrm{Co_{3}Sn_{2}S_{2}}$.
+(a) The conventional cell (black lines) and primitive cell (green lines) of $\mathrm{Co_{3}Sn_{2}S_{2}}$.
+(b) The magnetic phonon dispersion with phonon Zeeman splittings along $\Gamma$-T path.
+(c) The magnetic phonon dispersion of $E_{g}$ modes.
+The color gradient from red to blue represents the phonon angular momentum.
+Upper insets: The phonon magnetic moment for the two branches of $E_{g}$ mode along $\Gamma$-T path.
+Lower insets: Schematic of the two phonon modes with opposite angular momenta, corresponding to right-handed and left-handed circular motions of S atoms, respectively.
+Intrinsic phonon Chern states and the inertially decoupled model.
+The accurate calculation of phonon Zeeman splitting enables the determination of magnetic phonon spectra throughout the Brillouin zone, which is indispensable for exploring magnetic phonon topology.
+Following the analogy to the Chern state in the Haldane model [72], we adopt the two-dimensional (2D) honeycomb lattice [45, 73, 74] as an illustrative example.
+When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at KK($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $\Delta_{I}$ and $\Delta_{T}$, respectively.
+It has been shown that when $|\Delta_{T}|>|\Delta_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [45].
 Figure 3: Inertially decoupled model (IDM) and intrinsic phonon Chern states.
-(a) Schematic of chiral phonon modes in the high frequency regime: When $m\sim M$, both atoms contribute to high-frequency modes (upper panel); in the IDM condition ($m\ll M$), high-frequency modes are primarily contributed by light atoms (lower panel).
+(a) Schematic of circularly polarized phonon modes in the high frequency regime: When $m\sim M$, both atoms contribute to high-frequency modes (upper panel); in the IDM condition ($m\ll M$), high-frequency modes are primarily contributed by light atoms (lower panel).
 (b) Phonon dispersions of nonmagnetic honeycomb lattice and ferromagnetic triangular lattice.
@@
 Parameters are detailed in the SM [58].
-Intrinsic phonon Chern states and the inertially decoupled model.
-The accurate calculation of phonon Zeeman splitting enables the determination of magnetic phonon spectra throughout the Brillouin zone, which is indispensable for exploring magnetic phonon topology.
-Following the analogy to the Chern state in the Haldane model [48], we adopt the two-dimensional (2D) honeycomb lattice [49, 50, 51] as an illustrative example.
-When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $\Delta_{I}$ and $\Delta_{T}$, respectively.
-It has been shown that when $|\Delta_{T}|>|\Delta_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [49].
 However, it is challenging to realize such phonon Chern states in realistic materials with honeycomb lattices.
@@
 Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered.
-Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $\Delta_{I}$, which will make the condition $|\Delta_{T}|>|\Delta_{I}|$ difficult to satisfy.
 To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference.
 A prototypical example of this model is constructed by combining a honeycomb lattice composed of nonmagnetic light atoms of mass $m$ with a triangular lattice of heavy magnetic atoms of mass $M$ ($m\ll M$), as shown in Figs.
-2(b,c).
+3(b,c).
 Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions are mainly contributed by the light and heavy atoms, respectively, as shown in Figs.
-2(a,c), which is termed “inertially decoupled” (see SM [44] for details).
+3(a,c), which is termed “inertially decoupled” (see SM [58] for details).
 Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $\Delta_{I}$ mass at both the $K$ and $K^{\prime}$ valleys.
-Meanwhile, the magnetic $M$ atoms break TRS to generate large Zeeman splittings of chiral phonons via SPC, producing the desired large $\Delta_{T}$ mass.
+Meanwhile, the magnetic $M$ atoms break TRS to generate large phonon Zeeman splittings, producing the desired large $\Delta_{T}$ mass.
 The large $\Delta_{T}$ opens a topologically nontrivial phonon gap (seen in Fig.
-2(d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by topologically protected edge states crossing the phonon gap, as shown in Fig.
-2(f).
+3(d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by topologically protected edge states crossing the phonon gap, as shown in Fig.
+3(f).
 The inertially decoupled model is also applicable to three-dimensional (3D) magnets.
@@
 A representative phonon dispersion of such a 3D bulk model is shown in Fig.
-2(e).
+3(e).
 Notably, since each $m$ sublayer realizes a Chern state with $|C|=1$, the system exhibits a 3D phonon Chern state.
 This configuration hosts multiple unidirectional phonon edge states resulting from a total Chern number $|C|=N_{z}$, where $N_{z}$ is the number of $m$ sublayers.
-Figure 3: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$.
-(a,b) The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$.
-(c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the SPC, respectively.
-(e) The phonon dispersion of bulk $\mathrm{EuSi_{2}}$ under open-boundary conditions.
-(f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the SPC, respectively.
-(h) The phonon dispersion of the sandwiched slab under open-boundary conditions.
-The color gradient from red to blue represents the phonon angular momentum.
-Cyan and magenta colors in (c) and (h) label topological boundary modes residing on opposite boundaries.
-Candidate materials.
-Building upon our theory, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density-functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons.
-Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from the dynamical matrix through near-degenerate perturbation theory (see SM [44] for details), yielding TRS-breaking phonon dispersions.
-Through first-principles calculations guided by the inertially decoupled model, we identified five hexagonal ($P6/mmm$) magnets exhibiting intrinsic phonon Chern states, namely, $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [59], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [60], and $\mathrm{CrB_{2}}$ [61, 62].
+Through ab initio calculations guided by the inertially decoupled model, we identified five hexagonal ($P6/mmm$) magnets exhibiting intrinsic phonon Chern states, namely, $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ [75], $\mathrm{EuGa_{2}}$, $\mathrm{GdGa_{2}}$ [76], and $\mathrm{CrB_{2}}$ [77, 78].
 Remarkably, they all have large observable topological phonon gaps at the $\mathrm{K}$ point defined as $\Delta_{\text{ph}}=\Delta_{\text{z}}-\gamma$, where $\Delta_{\text{z}}=2|E_{\text{Z}}|$ is the phonon Zeeman splitting and $\gamma$ is the phonon linewidth resulting from EPC [79, 80], as summarized in Table I.
@@
 The lattice structures of $\mathrm{EuSi_{2}}$ are illustrated in Figs.
-3(a, b).
+4(a, b).
 Nonmagnetic $\mathrm{Si}$ atoms form a honeycomb-lattice layer, stacked alternately with a ferromagnetic $\mathrm{Eu}$ triangular-lattice layer, along the [001] direction.
-The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately $28\,u$ and $152\,u$ ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$.
+The atomic masses of $\mathrm{Si}$ and $\mathrm{Eu}$ are approximately 28u28\,u and 152u152\,u ($u$ is atomic mass unit), respectively, with each $\mathrm{Eu}$ atom carrying an out-of-plane magnetic moment of $7\,\mu_{B}$.
 As a result, $\mathrm{EuSi_{2}}$ can be aptly described by our proposed inertially decoupled model.
-Our first-principles calculations reveal that the EPC induces a pronounced Zeeman splitting of chiral phonons in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs.
-3(c, d)], opening a sizable full topological gap ($\approx 0.35$ THz) throughout the Brillouin zone.
+Figure 4: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$.
+(a,b) The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$.
+(c) The Brillouin zones of bulk $\mathrm{EuSi_{2}}$ and the projected side plane.
+(d,e) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (d) and with (e) the phonon Zeeman splittings, respectively.
+The calculated PMMs at $K(K^{\prime})$ and $H(H^{\prime})$ point are $5.7\times 10^{-3}\mu_{B}$ and $2.3\times 10^{-2}\mu_{B}$, respectively.
+(f) The phonon dispersion of bulk $\mathrm{EuSi_{2}}$ under open-boundary conditions.
+(g,h) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (g) and with (h) the phonon Zeeman splittings, respectively.
+The PMM at $K(K^{\prime})$ point is $2.2\times 10^{-2}\mu_{B}$.
+(i) The phonon dispersion of the sandwiched slab under open-boundary conditions.
+The color gradient from red to blue represents the phonon angular momentum.
+Cyan and magenta colors in (f) and (i) label topological boundary modes residing on opposite boundaries.
+Our ab initio calculations reveal that EPC-enhanced PMM ($\sim 10^{-2}\mu_{B}$) induces phonon Zeeman splitting in the high-frequency phonon dispersion of ferromagnetic $\mathrm{EuSi_{2}}$ [see Figs.
+4(d, e)], opening a sizable full topological gap ($\approx 0.35$ THz).
 Given that the Chern number remains $|C|=1$ for all 2D $k_{z}$ slices, $\mathrm{EuSi_{2}}$ realizes a 3D layered phonon Chern state.
 This is further supported by the presence of boundary modes in the phonon dispersions of $\mathrm{EuSi_{2}}$ under open-boundary conditions, as illustrated in Fig.
-3(e).
+4(f).
 To achieve a 2D Chern state, we construct a dynamically-stable sandwiched slab featuring a $\mathrm{Si}$ sublayer between two $\mathrm{Eu}$ sublayers, as shown by the black rectangle in Fig.
-3(b).
+4(b).
 By comparing Figs.
-2(f) and 2(g), we can see that the phonon Zeeman splitting generates a substantial topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ featuring a pair of edge modes traversing the phonon gap, as shown in Fig.
-3(h).
+4(g) and 4(h), we can also see that the phonon Zeeman splitting generates a topologically nontrivial phonon gap ($\approx 0.30$ THz) with a Chern number of $|C|=1$ featuring a pair of edge modes traversing the phonon gap, as shown in Fig.
+4(i).
 This provides definitive evidence for intrinsic phonon Chern states.
@@
 $\mathrm{CrB_{2}}$ 0.270 0.033 0.237
-Discussion and conclusion.
-The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring strong external magnetic fields, Coriolis forces, or stress gradients [2, 49].
-One potential advantage may be its ability to detect neutral particles by converting them into topological phonon edge currents.
-Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging.
-However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within the first “conduction” band above the topological gap in the phonon dispersion, corresponding phonon modes can be excited [65, 66], driving the system out of equilibrium.
-As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via topologically protected edge states, manifesting as a unidirectional edge phonon current immune to scattering [Fig.
-4(a)].
-This robust edge current is significantly more measurable than conventional delocalized bulk phonon excitations in materials.
-However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents.
-Thus, a large phonon Chern gap is highly desired to elevate the working temperature.
-Figure 4(b) lists phonon Chern gaps and transferred energy ranges of proposed candidate materials (see SM [44] for details), with transferred energy spanning approximately $20\sim 110\,\mathrm{meV}$.
-This energy range corresponds to low-energy dark matter which recently attracts much attention, therefore our work provides a potential detection mechanism for this category of dark matter candidates [65, 66].
-Figure 4: Detection of neutral particles based on the intrinsic phonon Chern state.
-(a) Schematic illustrations of the detection process.
-Left: Neutral particles (red beam) excite phonon modes (yellow circles in the lower panel) in high-frequency bands.
-Right: These phonon modes gradually relax to lower-frequency regimes via topologically protected edge phonon channels (orange and green lines across the gap), generating robust unidirectional phonon edge currents (orange and green arrows).
-(b) Transferred energy ranges and phonon Chern gaps from proposed candidate materials [Tab.
-S2 in SM [44]].
-In summary, we have established an ab initio theory enabling systematic predictions of EPC-induced magnetic properties of phonons and TRS-breaking phonon dispersions for insulators, semimetals and metals.
-Based on our theory, we construct the inertially decoupled model to seek intrinsic phonon Chern states in realistic materials, and propose candidates exhibiting EPC-induced large topological phonon gaps exceeding 0.3 THz.
-Furthermore, we demonstrate the potential capability of these topological phonon states to detect neutral particles through robust phonon edge currents.
-These advances could open broad pathways for investigating phonon magnetic effects in quantum materials.
+Conclusion.
+Based on the linear response theory, we have established an ab initio theory, enabling systematic investigation of EPC-induced magnetic phonon spectra and phonon magnetic properties for a wide range of insulators, semimetals, and metals.
+This theoretical method is validated through the excellent agreement with the recent experiment results of magnetic Weyl semimetal $\mathrm{Co_{3}Sn_{2}S_{2}}$, providing a quantitative explanation of the phonon Zeeman splitting induced by magnetic order.
+We also propose an inertially decoupled model to seek intrinsic phonon Chern states in realistic materials, and predict $\mathrm{EuSi_{2}}$ family materials as candidates exhibiting large topological phonon gaps exceeding 0.3 THz, which could be applied to detect neutral particles, such as dark matter particles [58].
+Therefore, our work paves a broad way for investigating phonon magnetic effects and phonon topology in quantum materials.
 Acknowledgements.
@@
 2024YFA1409100, No.
-2021YFA1400400 and No.
-2022YFA1403602), the Natural Science Foundation of Jiangsu Province (No.
-BK20220032, BK20233001, BK20243011), the Natural Science Foundation of China (No.
-12074181, No.
-12104217, No.
-12174182, and No.
+2021YFA1400400, No.
+2022YFA1403602, and No.
+2023YFA1407001), the Natural Science Foundation of Jiangsu Province (No.
+BK20253012, BK20252117, BK20233001, BK20243011, and BK20220032), the Natural Science Foundation of China (No.
+12534007 and No.
 92365203), and the e-Science Center of Collaborative Innovation Center of Advanced Microstructures.